# LODGE: Level-of-Detail Large-Scale Gaussian Splatting with Efficient Rendering

### v1

@@ 2 Related Work
Novel view synthesis has received great attention recently, mainly thanks to the advent of neural radiance fields (NeRFs) [21] and later 3D Gaussian Splatting (3DGS) [9], which offered an on par alternative with substantially faster rendering. Subsequently, 3DGS was further extended in many ways to handle antialiasing [35, 19], to offer better geometry [36, 7], or to accommodate different camera models [13, 33]. There has been a lot of effort on making 3D Gaussian Splatting representation smaller by compressing attributes of Gaussians [19, 2, 6, 23]. Similarly to our approach, this also reduces the required memory and makes rendering faster. Other approaches modify the rendering procedure to make rendering more efficient [20, 15, 11, 3, 34]. Finally, a lot of interest has been directed towards optimizing scene representation from images captured under different time-of-day, seasons, or exposure levels [12, 24, 14, 37]. We consider these approaches orthogonal to ours as they can also be applied on top of our method. Similarly, some works [20, 16, 24] focused on removing the number of Gaussians by proposing alternative densification and pruning strategies [20, 16], or by pruning Gaussians with little/no contribution to the rendering process [6, 2, 24] - similar to our approach. While these methods make rendering efficient for small-scale scenes, very large scenes are still not tractable to render in real-time without sacrificing quality.
@@ 3 Method
Our method aims to make large-scale 3D Gaussian Splatting reconstruction render fast even on mobile devices. To this end, we introduce a novel level-of-detail (LOD) representation and chunk-based caching scheme (see Fig. 2), leading to fast rendering whilst reducing the memory footprint.
@@
Upon splatting of the 3D Gaussians, the 2D screen is split into tiles of $16$ pixels each, with each 2D Gaussian being assigned to all (possibly many) tiles with which it overlaps. Subsequently, the Gaussians for each tile are then sorted and alpha blended in the rasterization process using Eq. 1.
@@
Level of Detail (LOD) representation. We propose to represent 3D scenes as multiple sets of Gaussians that correspond to different levels of detail as visualized in Fig. 2. As motivated in the introduction, distant regions of the scene are often represented by a large number of Gaussians that contribute little to the final rendering. Yet, these Gaussians still need to be evaluated by the renderer, causing excessive memory usage and computations. Individual pixels with a large number of visible Gaussians (Gaussians projected into them) can significantly slow down rendering due to how the pixels are processed in $16\times 16$ patches. The main goal is, therefore, to reduce the number of pixels with a large number of visible Gaussians, as shown in Fig. 3.
Therefore, we propose to represent faraway regions with less detailed sets of Gaussians and nearby regions with more detailed sets. To this end, we define multiple LOD levels: $\mathcal{G}^{(l)};L>l>0$, with $\mathcal{G}^{(0)}$ denoting the most detailed set, obtained by optimizing on the original set of images [9, 10]. We assume each LOD level $\mathcal{G}^{(l)}$ is constructed (as explained in the next section) such that a sufficient rendering quality is achieved when observed from a distance of at least $d_{l}$. When rendering the LOD from a given camera pose, we select a subset of Gaussians from each LOD level (see Fig. 2). We call this set the ‘active Gaussians’, and for camera center $\mathbf{c}$, it is defined as
@@
where $d_{0}=0$, $d_{L}=\infty$, and $\mu^{(l)}_{i}$ is the mean of Gaussian $g_{i}$ from level $l$. As shown in Fig. 3, LOD rendering reduces long tail of the per-pixel visible Gaussians distribution accelerating the rendering.
@@
To build the lower-detail representations $\mathcal{G}^{(l)}$ that are viewed at distances larger than $d_{l}$, we copy Gaussians from $\mathcal{G}^{(0)}$ and add a smoothing 3D filter (Eq. 3) for depth $d_{l}$. While adding a smoothing 3D filter alone does not directly lead to fewer Gaussians, a significant number of Gaussians will become redundant, reducing their contributions in alpha compositing. Eventually, using the importance score from [24], we iteratively prune unused Gaussians. Note that we always employ a few fine-tuning steps to correct for errors introduced by pruning Gaussians. For the fine-tuning, we use LOD rendering with levels up to the currently optimized LOD level. More details in the supp. mat.
@@
Selecting depth thresholds. An important question remains: how should the depth thresholds $d_{l}$ be selected to maximize the rendering performance. The renderer operates on pixels grouped in $16\times 16$ tiles and the rendering performance is heavily influenced by the number of Gaussians processed inside the same $16\times 16$-pixels tile. This is because all threads within a tile process the union of all visible Gaussians and must wait for the slowest thread to complete rasterizing. While it is difficult to theoretically analyze the impact of thresholds on the number of processed Gaussians, we can estimate this number (cost) for different thresholds - by rendering a subset of training views - and choose the one that minimizes the average number of Gaussians per tile. In Fig. 4, we show the cost distribution for two LOD levels at depths $d_{1}$ and $d_{2}$. Notice how the value of the minimum is similar for set $\{(x,ax+b):x\in\mathbb{R}\}$. This enables us to leverage a simple greedy strategy - starting at $d_{1}$, iteratively adding more thresholds. This reduces the complexity of the search problem to a linear one.
@@
Reducing memory with chunk-based rendering. While LOD rendering (described so far) makes rasterization fast by reducing the number of visible Gaussians, all Gaussians still need to be loaded in GPU memory. Moreover, the ‘active Gaussians’ need to be continuously recomputed, leading to overhead computations. This can be prohibitive for small devices where the amount of available memory is limited. To address this, we propose to split the scene into different regions called ‘chunks’ and store a fixed set of ‘active Gaussians’ per chunk. When rendering an image, the rasterizer simply uses the ‘active Gaussians’ pre-computed for the closest chunk. We split the scene into chunks by performing a K-means clustering on the training camera positions (more details in supp. mat.). For each chunk with center $\mathbf{c}$, the ‘active Gaussians’ are defined by Eq. 2 computed at the chunk center (rather than the camera position), with depths $d_{l}$ offset by the chunk radius (distance to next closest chunk center) to ensure sufficient resolution for all camera positions inside the chunk.
@@
Visibility filtering. Given our LOD chunks, we further accelerate rendering by filtering the set of Gaussians for each LOD chunk. Following RadSplat [24] we additionally perform importance pruning for each LOD chunk, ie., we compute per-Gaussian importance scores and remove all Gaussians with importance score lower than a fixed threshold. However, to further increase robustness, we add additional views by adding random perturbations to existing training views within the LOD chunk. To this end, we employ the original camera positions, but sample random orientations.
@@
Opacity blending for smooth cross-chunk transitions. Having constructed these chunks, a naive approach would be to simply render the ‘active Gaussians’ for the chunk center. While this speeds up rendering and reduces memory, it also introduces sharp changes when rendering a dynamic video of a camera trajectory (see Fig. 7). These artifacts are inherently caused by an abrupt change in the ‘active Gaussians’ while moving, without employing any form of smoothing filter during the transition. To resolve the issue, we propose a smoothing strategy leveraging the two closest chunks (see Fig. 2). When rendering an image, we first seek the two closest chunk centers. We take their respective sets of active Gaussians and combine them. Afterwards, we modulate the opacity of Gaussians, which are not in the intersection of the two sets of active Gaussians, using
@@
where $\mathbf{c}$ is the current camera position, $\mathbf{m}_{f}$ is the chunk center to which Gaussian $i$ belongs, and $\mathbf{m}_{o}$ is the other chunk (not containing Gaussian i). Note that $\bar{t}$ is the normalized length of the projection of $(\mathbf{c}-\mathbf{m_{o}})$ onto the line connecting the two chunk centers. We use the length of projection instead of the Euclidean distance to achieve a smooth transition even when the camera doesn’t pass through the chunk center. Given the union of the two sets of active Gaussians with modified opacity, we proceed with the standard rasterization step. However, note that only the union of the two chunks needs to be loaded in memory and at each rendering pass, we only need to update the opacity of the symmetric difference of the two active Gaussian sets. In practice, reloading of LOD splits can be done in a background process that does not affect the renderer’s runtime.
@@ 4 Experiments
To validate our method, we conduct experiments on large-scale indoor and outdoor datasets, comparing our method to other SOTA approaches. We further analyze individual components of our method and evaluate rendering speed on various mobile devices. We report standard PSNR, SSIM, and LPIPS (VGG) metrics, but also FPS and the number of Gaussians loaded in GPU memory (#G). For more details on implementation and the evaluation protocol, we kindly refer the reader to the supp. mat. All baselines and our method use a single NVIDIA A100 SXM4 40GB GPU for training and evaluation. For mobile experiments, we used two iPhones (13 Mini and 15 Pro), as well as two lower-end laptops without a powerful GPU (MacBook Air M3 and HP Chromebook).
@@
Datasets & baselines. We use two larger-scale datasets to validate our approach, ie., two outdoor scenes from Hierarchical 3DGS dataset [10] and three indoor scenes from Zip-NeRF dataset [1]. Each scene consists of around 1000-2000 images. We compare our approach against the following baselines: Current SOTA on the Zip-NeRF dataset - Zip-NeRF [1] - slow to train and without real-time rendering. Traditional 3DGS baselines - 3DGS [9], Mip-Splatting [35], and Scaffold-GS [19]. LOD-based methods - H3DGS [10] (SOTA on the Hierarchical 3DGS dataset), FLOD [28], and Octree-GS [27]. Note that H3DGS, Octree-GS, FLOD were trained for 45K, 40K, and 100K iterations, while ours was trained for 36K iterations. For FLOD we use 3-4-5 LOD rendering and for H3DGS we employ their default $\tau=6$. More details can be found in the supp. mat.
Evaluation on Hierarchical 3DGS dataset. As shown in Tab. 1, Zip-NeRF struggles on these large scenes, despite requiring much longer training (200K iterations vs. 36K for ours). This likely stems from its use of a contiguous scene representation with limited resolution, unlike 3DGS, which sparsely encodes only occupied space. Among non-LOD baselines, 3DGS [9] and Mip-Splatting [35] offer competitive quality and rendering speed, whereas Scaffold-GS [19] lags behind, likely due to suboptimal densification in large-scale scenes. H3DGS [10] attains high visual fidelity, but incurs significant overhead from per-frame LOD splitting and the large number of Gaussians, making it slower than other methods. FLOD fails to generate sufficient Gaussians due to their coarse-to-fine training scheme, discarding fine details early and struggling to recover them later, even though it uses three times more training iterations than related methods. For Octree-GS [27], ours achieves higher overall quality. Interestingly, Octree-GS excels on Campus in PSNR but underperforms in SSIM and LPIPS. We attribute this to its MLP-based design, which appears more robust to exposure variation but less effective in preserving perceptual similarity. In summary, our method consistently achieves the best results across all scenes whilst maintaining superior rendering efficiency. In qualitative comparisons (Fig. 5), we highlight a close-up and distant regions to assess both local details and long-range fidelity. Overall, our method delivers sharper reconstructions in both close and distant regions. In contrast, Octree-GS appears blurrier nearby and suffers from color shifts at a distance, while H3DGS often exhibits jagged geometry, potentially due to its reliance on depth supervision.
@@
Performance analysis & design validation. We conduct an ablation study on the SmallCity scene to assess the impact of key components of our method on rendering quality and speed. Quantitative results are shown in Tab. 3, reporting PSNR, render time, number of visible Gaussians, and memory usage. Time for reloading Gaussians at chunk boundaries is excluded, as it is handled asynchronously. Qualitative results are shown in Fig. 7. We start with the base method (‘Full representation’), which includes importance pruning but no LOD. Adding up to four LOD levels significantly improves rendering speed (up to 3×3$\times$3 × with just one level) at a minimal cost in PSNR and sharpness due to more aggressive pruning. As additional levels yield diminishing returns, we decided to always employ two LODs for all experiments, as a good speed and quality trade-off. Thresholds were selected automatically (Section 3, ‘Selecting depth thresholds’), and we show that when instead setting them manually (15 and 70 m), performance worsens compared to our automatic selection ‘LOD (d=10,28)’.
The last three rows evaluate our chunk-based rendering. Clustering camera positions and adjusting depth thresholds by chunk radius improves quality (‘LOD (d=10,28) + clusters’), but increases Gaussian count due to finer LOD resolutions. Visibility filtering (‘+ vis. filtering’) further reduces both visible and loaded Gaussians, and thus rendering time. However, using the active Gaussians from only the chunk centers introduces visual artifacts at chunk boundaries, noticeable as sharpness discontinuities (Fig. 7, second column from right). To circumvent this issue, we introduce opacity blending between adjacent chunks (last row), which smoothens transitions. While it slightly increases render time and Gaussian count, it runs faster and more memory-efficient than its non-chunk-based LOD counterpart, being particularly beneficial for memory-constrained mobile devices.
@@
Rendering on mobile and low-power devices. Finally, we benchmark rendering speed on four devices: iPhone 13 Mini, iPhone 15 Pro, HP Elite Dragonfly Chromebook, and MacBook Air 13 inch, using the web-based 3DGS renderer by Mark Kellogg [8], which performs rendering with asynchronous Gaussian sorting running in parallel with the rasterization. We report both FPS (rasterization only) and sorting time (S. time) in milliseconds in Tab. 4. For H3DGS [10], we show both the full model ($\tau=0$) and the default ($\tau=6$) configuration. Vanilla 3DGS renders fast due to its low Gaussian count but at the cost of poor quality (see Tab. 1; PSNR @ 25.42 vs. 26.57 for ours). H3DGS instead fails on iPhones due to memory limitations and cannot render in real time on laptops. Notably, iPhone 13 Mini outperforms the 15 Pro in FPS due to its smaller display. Our method runs efficiently across all devices and achieves the best performance on laptops.

### v2

@@ 2 Related Work
Novel view synthesis has received great attention recently, mainly thanks to the advent of neural radiance fields (NeRFs) [22] and later 3D Gaussian Splatting (3DGS) [10], which offered an on par alternative with substantially faster rendering. Subsequently, 3DGS was further extended in many ways to handle antialiasing [37, 20, 31], to offer better geometry [38, 8], or to accommodate different camera models [14, 35]. There has been a lot of effort on making 3D Gaussian Splatting representation smaller by compressing attributes of Gaussians [20, 3, 7, 24]. Similarly to our approach, this also reduces the required memory and makes rendering faster. Other approaches modify the rendering procedure to make rendering more efficient [21, 16, 12, 4, 36, 42]. Finally, a lot of interest has been directed towards optimizing scene representation from images captured under different time-of-day, seasons, or exposure levels [13, 25, 15, 39]. We consider these approaches orthogonal to ours as they can also be applied on top of our method. Similarly, some works [21, 17, 25] focused on removing the number of Gaussians by proposing alternative densification and pruning strategies [21, 17], or by pruning Gaussians with little/no contribution to the rendering process [7, 3, 25] - similar to our approach. While these methods make rendering efficient for small-scale scenes, very large scenes are still not tractable to render in real-time without sacrificing quality.
@@ 3 Method
Our method aims to make large-scale 3D Gaussian Splatting reconstruction render fast even on mobile devices. To this end, we introduce a novel level-of-detail (LOD) representation and chunk-based caching scheme (see Fig.˜2), leading to fast rendering whilst reducing the memory footprint.
@@
Upon splatting of the 3D Gaussians, the 2D screen is split into tiles of $16$ pixels each, with each 2D Gaussian being assigned to all (possibly many) tiles with which it overlaps. Subsequently, the Gaussians for each tile are then sorted and alpha blended in the rasterization process using Eq.˜1.
@@
Level of Detail (LOD) representation. We propose to represent 3D scenes as multiple sets of Gaussians that correspond to different levels of detail as visualized in Fig.˜2. As motivated in the introduction, distant regions of the scene are often represented by a large number of Gaussians that contribute little to the final rendering. Yet, these Gaussians still need to be evaluated by the renderer, causing excessive memory usage and computations. Individual pixels with a large number of visible Gaussians (Gaussians projected into them) can significantly slow down rendering due to how the pixels are processed in $16\times 16$ patches. The main goal is, therefore, to reduce the number of pixels with a large number of visible Gaussians, as shown in Fig.˜3.
Therefore, we propose to represent faraway regions with less detailed sets of Gaussians and nearby regions with more detailed sets. To this end, we define multiple LOD levels: $\mathcal{G}^{(l)};L>l>0$, with $\mathcal{G}^{(0)}$ denoting the most detailed set, obtained by optimizing on the original set of images [10, 11]. We assume each LOD level $\mathcal{G}^{(l)}$ is constructed (as explained in the next section) such that a sufficient rendering quality is achieved when observed from a distance of at least $d_{l}$. When rendering the LOD from a given camera pose, we select a subset of Gaussians from each LOD level (see Fig.˜2). We call this set the ‘active Gaussians’, and for camera center $\mathbf{c}$, it is defined as
@@
where $d_{0}=0$, $d_{L}=\infty$, and $\mu^{(l)}_{i}$ is the mean of Gaussian $g_{i}$ from level $l$. As shown in Fig.˜3, LOD rendering reduces long tail of the per-pixel visible Gaussians distribution accelerating the rendering.
@@
To build the lower-detail representations $\mathcal{G}^{(l)}$ that are viewed at distances larger than $d_{l}$, we copy Gaussians from $\mathcal{G}^{(0)}$ and add a smoothing 3D filter (Eq.˜3) for depth $d_{l}$. While adding a smoothing 3D filter alone does not directly lead to fewer Gaussians, a significant number of Gaussians will become redundant, reducing their contributions in alpha compositing. Eventually, using the importance score from [25], we iteratively prune unused Gaussians. Note that we always employ a few fine-tuning steps to correct for errors introduced by pruning Gaussians. For the fine-tuning, we use LOD rendering with levels up to the currently optimized LOD level. More details in the supp. mat.
@@
Selecting depth thresholds. An important question remains: how should the depth thresholds $d_{l}$ be selected to maximize the rendering performance. The renderer operates on pixels grouped in $16\times 16$ tiles and the rendering performance is heavily influenced by the number of Gaussians processed inside the same $16\times 16$-pixels tile. This is because all threads within a tile process the union of all visible Gaussians and must wait for the slowest thread to complete rasterizing. While it is difficult to theoretically analyze the impact of thresholds on the number of processed Gaussians, we can estimate this number (cost) for different thresholds - by rendering a subset of training views - and choose the one that minimizes the average number of Gaussians per tile. In Fig.˜4, we show the cost distribution for two LOD levels at depths $d_{1}$ and $d_{2}$. Notice how the value of the minimum is similar for set $\{(x,ax+b):x\in\mathbb{R}\}$. This enables us to leverage a simple greedy strategy - starting at $d_{1}$, iteratively adding more thresholds. This reduces the complexity of the search problem to a linear one.
@@
Reducing memory with chunk-based rendering. While LOD rendering (described so far) makes rasterization fast by reducing the number of visible Gaussians, all Gaussians still need to be loaded in GPU memory. Moreover, the ‘active Gaussians’ need to be continuously recomputed, leading to overhead computations. This can be prohibitive for small devices where the amount of available memory is limited. To address this, we propose to split the scene into different regions called ‘chunks’ and store a fixed set of ‘active Gaussians’ per chunk. It is important to note that the set of ‘active Gaussians’ associated with a chunk is not merely a subset of all Gaussians located within its spatial boundaries (see Fig.˜2. Instead, it represents the entire scene, with regions closer to the chunk center modeled at higher levels of detail. When rendering an image, the rasterizer simply uses the ‘active Gaussians’ pre-computed for the closest chunk.
To define the chunks, we split the scene by performing a K-means clustering on the training camera positions (more details in supp. mat.). For each chunk with center $\mathbf{c}$, the ‘active Gaussians’ are defined by Eq.˜2 computed at the chunk center (rather than the camera position), with depths $d_{l}$ offset by the chunk radius (distance to next closest chunk center) to ensure sufficient resolution for all camera positions inside the chunk.
@@
Visibility filtering. Given our LOD chunks, we further accelerate rendering by filtering the set of Gaussians for each LOD chunk. Following RadSplat [25] we additionally perform importance pruning for each LOD chunk, ie., we compute per-Gaussian importance scores and remove all Gaussians with importance score lower than a fixed threshold. However, to further increase robustness, we add additional views by adding random perturbations to existing training views within the LOD chunk. To this end, we employ the original camera positions, but sample random orientations. Note, that if orientations were included alongside positions in defining LOD chunks, visibility filtering could achieve an additional reduction in the number of loaded Gaussians. Nonetheless, this gain would not be significant and would reduce the method’s ability to handle rapid camera rotations and wide fields of view, both of which are common in practical applications (e.g. AR/VR).
@@
Opacity blending for smooth cross-chunk transitions. Having constructed these chunks, a naive approach would be to simply render the ‘active Gaussians’ for the chunk center. While this speeds up rendering and reduces memory, it also introduces sharp changes when rendering a dynamic video of a camera trajectory (see Fig.˜7). These artifacts are inherently caused by an abrupt change in the ‘active Gaussians’ while moving, without employing any form of smoothing filter during the transition. To resolve the issue, we propose a smoothing strategy leveraging the two closest chunks (see Fig.˜2). When rendering an image, we first seek the two closest chunk centers. We take their respective sets of active Gaussians and combine them. Afterwards, we modulate the opacity of Gaussians, which are not in the intersection of the two sets of active Gaussians, using
@@
where $\mathbf{c}$ is the current camera position, $\mathbf{m}_{f}$ and $\mathbf{m}_{o}$ are the two closest chunk centers – $\mathbf{m}_{f}$ being the chunk center to which Gaussian $i$ belongs, and $\mathbf{m}_{o}$ being the other chunk (not containing Gaussian i). Note that $\bar{t}$ is the normalized length of the projection of $(\mathbf{c}-\mathbf{m_{o}})$ onto the line connecting the two chunk centers. We use the length of projection instead of the Euclidean distance to achieve a smooth transition even when the camera doesn’t pass through the chunk center. Given the union of the two sets of active Gaussians with modified opacity, we proceed with the standard rasterization step. However, note that only the union of the two chunks needs to be loaded in memory and at each rendering pass, we only need to update the opacity of the symmetric difference of the two active Gaussian sets.
In practice, reloading of LOD splits can be done in a background process that does not affect the renderer’s runtime. We begin by loading the active Gaussians’ properties (e.g., positions, colors) into GPU memory. As the camera moves, we only need to modulate the opacity of the two loaded chunks until the camera passes the center of the chunk. Then, we remove the Gaussians from the previous chunk, keep the ones from the closest one, and load Gaussians from the second closest chunk. The opacity blending at this point assigns weights close to 1 to all Gaussians from the closest chunk so there are no artifacts during loading the next chunk – even if loading of the next chunk is delayed.
@@ 4 Experiments
To validate our method, we conduct experiments on large-scale indoor and outdoor datasets, comparing our method to other SOTA approaches. We further analyze individual components of our method and evaluate rendering speed on various mobile devices. We report standard PSNR, SSIM, and LPIPS (VGG) metrics, but also FPS and the number of Gaussians loaded in GPU memory (#G). The number of loaded Gaussians serves as a proxy for memory usage, offering a fairer comparison across 3DGS implementations (as opposed to peak-memory usage) as it does not depend on the actual implementations but rather on the algorithm itself. For more details on implementation and the evaluation protocol, we kindly refer the reader to the supp. mat. All baselines and our method use a single NVIDIA A100 SXM4 40GB GPU for training and evaluation. For mobile experiments, we used two iPhones (13 Mini and 15 Pro), as well as two lower-end laptops without a powerful GPU (MacBook Air M3 and HP Chromebook).
@@
Datasets & baselines. We use two larger-scale datasets to validate our approach, ie., two outdoor scenes from Hierarchical 3DGS dataset [11] and three indoor scenes from Zip-NeRF dataset [2]. Each scene consists of around 1000-2000 images. We compare our approach against the following baselines: Current SOTA on the Zip-NeRF dataset - Zip-NeRF [2] - slow to train and without real-time rendering. Traditional 3DGS baselines - 3DGS [10], Mip-Splatting [37], and Scaffold-GS [20]. LOD-based methods - H3DGS [11] (SOTA on the Hierarchical 3DGS dataset), FLOD [29], Octree-GS [28], and CityGS [18]. Note that H3DGS, Octree-GS, FLOD were trained for 45K, 40K, and 100K iterations, while ours was trained for 36K iterations. For FLOD we use 3-4-5 LOD rendering and for H3DGS we employ their default $\tau=6$. More details can be found in the supp. mat.
Evaluation on Hierarchical 3DGS dataset. As shown in Tab.˜1, Zip-NeRF struggles on these large scenes, despite requiring much longer training (200K iterations vs. 36K for ours). This likely stems from its use of a contiguous scene representation with limited resolution, unlike 3DGS, which sparsely encodes only occupied space. Among non-LOD baselines, 3DGS [10] and Mip-Splatting [37] offer competitive quality and rendering speed, whereas Scaffold-GS [20] lags behind, likely due to suboptimal densification in large-scale scenes. H3DGS [11] attains high visual fidelity, but incurs significant overhead from per-frame LOD splitting and the large number of Gaussians, making it slower than other methods. FLOD fails to generate sufficient Gaussians due to their coarse-to-fine training scheme, discarding fine details early and struggling to recover them later, even though it uses three times more training iterations than related methods. Compared to Octree-GS [28] and CityGS [18], ours achieves higher overall quality while being twice as fast. Interestingly, both Octree-GS and CityGS excel on Campus in PSNR but underperform in SSIM and LPIPS. We attribute this to our method being more robust to exposure variation but less effective in preserving perceptual similarity under appearance changes. In summary, our method consistently achieves the best results across all scenes whilst maintaining superior rendering efficiency. In qualitative comparisons (Fig.˜5), we highlight a close-up and distant regions to assess both local details and long-range fidelity. Overall, our method delivers sharper reconstructions in both close and distant regions. In contrast, Octree-GS appears blurrier nearby and suffers from color shifts at a distance, while H3DGS often exhibits jagged geometry, potentially due to its reliance on depth supervision.
@@
In the Hierarchical 3DGS dataset [11], the camera follows a linear trajectory with chunks aligned along it; therefore, only the two nearest chunks are relevant. In Zip-NeRF [2], however, the chunks fully cover the scene and the camera may traverse diagonally, potentially being between three or more chunks – raising the question of whether opacity blending of only two chunks is sufficient. Empirically, no artifacts were observed across Hierarchical 3DGS [11], Zip-NeRF [2], and Mip-NeRF 360 [1]. We hypothesize that: a) Close to the training camera distribution, neighboring chunks are well constrained, producing near-identical renderings even without blending. b) Far from the distribution, viewpoints lie at chunk boundaries where only the two closest chunks dominate. Thus, two-chunk blending is generally sufficient.
@@
Performance analysis & design validation. We conduct an ablation study on the SmallCity scene to assess the impact of key components of our method on rendering quality and speed. Quantitative results are shown in Tab.˜3 (corresponding to SmallCity in Tab.˜1), reporting PSNR, render time, number of visible Gaussians, and memory usage. Time for reloading Gaussians at chunk boundaries is excluded, as it is handled asynchronously. Qualitative results are shown in Fig.˜7. We start with the base method (‘Full representation’), which includes importance pruning but no LOD. Adding up to four LOD levels significantly improves rendering speed (up to $3\times$ with just one level) at a minimal cost in PSNR and sharpness due to more aggressive pruning. As additional levels yield diminishing returns, we decided to always employ two LODs for all experiments, as a good speed and quality trade-off. Thresholds were selected automatically (Section˜3, ‘Selecting depth thresholds’), and we show that when instead setting them manually (15 and 70 m), performance worsens compared to our automatic selection ‘LOD (d=10,28)’.
The last three rows evaluate our chunk-based rendering. Clustering camera positions and adjusting depth thresholds by chunk radius improves quality (‘LOD (d=10,28) + clusters’), but increases Gaussian count due to finer LOD resolutions. Visibility filtering (‘+ vis. filtering’) further reduces both visible and loaded Gaussians, and thus rendering time. However, using the active Gaussians from only the chunk centers introduces visual artifacts at chunk boundaries, noticeable as sharpness discontinuities (Fig.˜7, second column from right). To circumvent this issue, we introduce opacity blending between adjacent chunks (last row, corresponding to ‘Ours’ in Tab.˜1), which smoothens transitions. While it slightly increases render time and Gaussian count, it runs faster and more memory-efficient than its non-chunk-based LOD counterpart, being particularly beneficial for memory-constrained mobile devices.
@@
Rendering on mobile and low-power devices. Finally, we benchmark rendering speed on four devices: iPhone 13 Mini, iPhone 15 Pro, HP Elite Dragonfly Chromebook, and MacBook Air 13 inch, using the web-based 3DGS renderer by Mark Kellogg [9], which performs rendering with asynchronous Gaussian sorting running in parallel with the rasterization. We report both FPS (rasterization only) and sorting time (S. time) in milliseconds in Tab.˜4. For H3DGS [11], we show both the full model ($\tau=0$) and the default ($\tau=6$) configuration. Vanilla 3DGS renders fast due to its low Gaussian count but at the cost of poor quality (see Tab.˜1; PSNR @ 25.42 vs. 26.57 for ours). H3DGS instead fails on iPhones due to memory limitations and cannot render in real time on laptops. Notably, iPhone 13 Mini outperforms the 15 Pro in FPS due to its smaller display. Our method runs efficiently across all devices and achieves the best performance on laptops.